An Airplane climbs at an angle of 11 degrees at an average speed of 420 mph. How long will it take for the plane to reach its cruising altitude of 6.5 miles? Round to nearest minute.
Answer by rothauserc(4718)

let x be the value of the hypotenuse
:
sin(11) = 6.5 / h
:
h = 6.5 / sin(11) = 6.5 / 0.1908 = 34.0655
:
rate * time = distance
:
time = distance / rate = 34.0655 / 420 = 0.0811
:
convert time to minutes
:
*********************************************
0.0811 * 60 = 4.866 approximately 5 minutes
*********************************************
: